OpenStudy (toxicsugar22):

Duncan knows that it takes 36400cal of energy to heat a pint of water from room temperature to boiling. However, Duncan has prepared ramen noodles so many times he does not need to measure the water carefully. If he happens to heat 0.600pint of room-temperature water, how many kilojoules of heat energy will have been absorbed by the water at the moment it begins to boil? Express your answer numerically in kilojoules.

OpenStudy (anonymous):

Hmm, looks like a proportion. The new amount of energy is 0.600 times 36.400 kJ/pint

OpenStudy (toxicsugar22):

21.84

OpenStudy (anonymous):

Unless I'm missing something, I think it is just that simple. Though, for the sake of significant figures, I'd round to 21.8 kJ

OpenStudy (toxicsugar22):

do you have any idea carlos

OpenStudy (anonymous):

Oh, woops, forgot to convert calories to joules... Derp

OpenStudy (anonymous):

Sorry, I was missing something.

OpenStudy (toxicsugar22):

so 21.8 is the calories and we convert to kilojoules

OpenStudy (anonymous):

21.84 is kilocalories because already divided by 1000

OpenStudy (anonymous):

It is not right. If you need 36400 cal for 1 pint, for 0.6 pints yo need 0.6x36400=21840 cal=21.840 Kcal Now you need to convert Kcal to Kjoules: 1Kcal=4.1868 KJoules, then the solution is: 91.44 KJoules

OpenStudy (toxicsugar22):

yes that is what i got

OpenStudy (toxicsugar22):

thansk carlos

OpenStudy (anonymous):

welcome

OpenStudy (toxicsugar22):

can you do one more

OpenStudy (anonymous):

shoot

OpenStudy (toxicsugar22):

water (3110g ) is heated until it just begins to boil. If the water absorbs 5.01×105J of heat in the process, what was the initial temperature of the water?

OpenStudy (toxicsugar22):

i dont get this one at all

OpenStudy (toxicsugar22):

do you get it

OpenStudy (anonymous):

\[Q=m·c·\Delta T\] We know it is heated until it boils, then \[\Delta T=100-T_0\] and \[c=4.18 KJ·Kg^{-1}·ºC^{-1}\]\[100-T_0=\frac{ 5.01·10^2 }{ 4.18·3.11 }=38.5 \rightarrow T_0=100-38.5=61.5\]in Celsius Note that I have converted gr to Kgr and Joules to KJoules to match the constant "c" units

OpenStudy (anonymous):

Note also that by definition: 1Kcal increases the temperature of 1 liter of water in 1 Degree Celsius. Converting 1 Kcal = 4.18KJ you get the constant I have given you

OpenStudy (anonymous):

@toxicsugar22 solution is ready

OpenStudy (toxicsugar22):

ok

OpenStudy (toxicsugar22):

so 61.5 is the solution

OpenStudy (toxicsugar22):

you made theis problem so easy

OpenStudy (anonymous):

Can you assess the solution?

OpenStudy (toxicsugar22):

yes so is 61.5 the answer

OpenStudy (anonymous):

so we got it right. Its been more than 30 years I have not solved problems of Thermodynamics

OpenStudy (toxicsugar22):

thanks you you are the best tutor ever

OpenStudy (toxicsugar22):

you explain everthing so simply

OpenStudy (anonymous):

not at all. There are very good people around here, really knowledgeable

OpenStudy (toxicsugar22):

UP NOPE IT WAS NOT RIGHT

OpenStudy (anonymous):

The problem was easy itself. Sometimes I become entangled in my own explanations

OpenStudy (toxicsugar22):

i put it in my homeowork

OpenStudy (toxicsugar22):

and it says it not right

OpenStudy (anonymous):

Hoe many decimal places youve got to give?

OpenStudy (toxicsugar22):

it says Enter your answer using dimensions of temperature

OpenStudy (toxicsugar22):

decimals are fine

OpenStudy (anonymous):

I see no other solution: 1 calory increases 1 cc of water (1g) by one degree celsius. 1 Kcal increase 1C in one liter. The rest is easy

OpenStudy (toxicsugar22):

like how so was it right 61.5

OpenStudy (toxicsugar22):

and then we hoave to add hte units

OpenStudy (anonymous):

Exat solution is: 61.4609456 C that I have rounded up to 1 decimal place 61.5

OpenStudy (toxicsugar22):

ok and the units

OpenStudy (anonymous):

Unites are degrees celsius

OpenStudy (anonymous):

Maybe you have to put 61.5C

OpenStudy (toxicsugar22):

i did

OpenStudy (anonymous):

and did not work?

OpenStudy (toxicsugar22):

NO

OpenStudy (toxicsugar22):

MAYBE IT IS RIGHT BUT I DONT KNOW WHAY THEY ARE COUNTING IT AS WRONG

OpenStudy (toxicsugar22):

NO 61.5 WAS CORECT SORRY

OpenStudy (toxicsugar22):

FORGOT TO PUT DEGREE SIGN

OpenStudy (anonymous):

oops!